A cord 10m long is hung between two posts 6m apart. What maximum load hung at the middle can the system support if the tensile strength of the cord is 350N?


1
Expert's answer
2021-12-15T16:21:51-0500

"2T\cos\alpha=mg\to m=2T\cos\alpha/g="


"=2\cdot350\cdot(4/5)/9.8=57\ (kg)"